OpenStudy (anonymous):

what is 10z/ squre root of 3

OpenStudy (anonymous):

\[\frac{ 10z }{ \sqrt{3}} = \frac{ 10z }{ \sqrt{3}} * \frac{\sqrt{3}}{ \sqrt{3}} \\ \frac{10z \sqrt{3}}{3} \]

OpenStudy (anonymous):

Thanks your my hero ^_^ lol